OpenStudy (anonymous):

CONFUSED! PLEASEE HELP!? Zach borrowed 4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved 5,920 to pay back to his grandfather. At what interest rate did Zach borrow the money?

jimthompson5910 (jim_thompson5910):

A = P + Prt 5920 = 4000 + 4000*r*3 5920 - 4000 = 4000*r*3 1920 = 12000*r 1920/12000 = r 0.16 = r r = 0.16 So the interest rate was 16%